# Arcade Processes for Informed Martingale Interpolation

### v2

@@ 4 Filtered arcade martingales
    $\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n}):=\{\pi\in\Pi(\mu_{0},\mu_{1},\ldots,\mu_{n})\cap\mathcal{P}_{1}(\mathbb{R}^{n+1})\,|\,(X_{0},X_{1},\ldots,X_{n})\sim\pi\implies\\ \mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m}]=X_{m}\text{ for all }m=0,\ldots,n\}$ is the set of martingale couplings of $\{\mu_{0},\mu_{1},\ldots,\mu_{n}\}\subseteq\mathcal{P}(\mathbb{R})$.
We recall that $\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})\neq\emptyset$ if and only if $\int_{\mathbb{R}}x\,\mathrm{d}\mu_{0}(x)=\int_{\mathbb{R}}x\,\mathrm{d}\mu_{1}(x)=\ldots=\int_{\mathbb{R}}x\,\mathrm{d}\mu_{n}(x)$ and $\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{0}(x)\leqslant\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{1}(x)\leqslant\ldots\leqslant\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{n}(x)$ for any convex function $f$ on $\mathbb{R}$, i.e, the measures $\mu_{0},\mu_{1},\ldots,\mu_{n}$ are convexly ordered. In that case, we write $\mu_{0}\leqslant_{\mathrm{cx}}\mu_{1}\leqslant_{\mathrm{cx}}\ldots\leqslant_{\mathrm{cx}}\mu_{n}$.
@@ 4.1 The one-arc FAM
Let $X=(X_{0},X_{1})$ be a random vector distributed according to a martingale coupling $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1})$, where $(\mu_{0},\mu_{1})\in\mathcal{P}_{1}(\mathbb{R})\times\mathcal{P}_{1}(\mathbb{R})$ and $\mu_{0}\leqslant_{\mathrm{cx}}\mu_{1}$. It follows that $X_{0}\sim\mu_{0}$, $X_{1}\sim\mu_{1}$, and $\mathbb{E}[X_{1}\,|\,X_{0}]=X_{0}$. Given an $X$-RAP $(I_{t}^{(1)})_{t\in[T_{0},T_{1}]}$ on the partition $\{T_{0},T_{1}\}_{*}$, we aim at constructing an $(\mathcal{F}^{I}_{t})$-martingale $(M_{t})_{t\in[T_{0},T_{1}]}$ such that $M_{T_{0}}\overset{a.s.}{=}X_{0}$ and $M_{T_{1}}\overset{a.s.}{=}X_{1}$. The BHM framework developed in [7] is recovered when $X_{0}=\mathbb{E}[X_{1}]=0$ and $T_{0}=0$.
@@
Given an $X$-RAP $(I_{t}^{(1)})$, a one-arc FAM for $X\sim\pi^{X}$ on $[T_{0},T_{1}]$ is a stochastic process defined $M_{t}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{t}^{I}]$.
@@
    For $s<t$, $\mathbb{E}[M_{t}\,|\,\mathcal{F}_{s}^{I}]=M_{s}$ by the tower property of the conditional expectation.
@@
    $M_{T_{0}}=\mathbb{E}[X_{1}\,|\,X_{0}]\overset{a.s.}{=}X_{0}$, since $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1})$ and $I_{T_{0}}^{(1)}=X_{0}$.
@@
    $M_{T_{1}}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{T_{1}}^{I}]\overset{a.s.}{=}X_{1}$, since $\sigma(X_{1})\subset\mathcal{F}_{T_{1}}^{I}$ by construction of $(I_{t}^{(1)})$.
Hence, $(M_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale that interpolates almost surely between $X_{0}$ and $X_{1}$ over the interval $[T_{0},T_{1}]$. ∎
@@
$\mathbb{E}[M_{t}\,|\,\mathcal{F}^{M}_{s}]=\mathbb{E}[\mathbb{E}[M_{t}\,|\,\mathcal{F}^{M}_{s}]\,|\,\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[M_{t}\,|\,\mathcal{F}^{I}_{s}]\,|\,\mathcal{F}^{M}_{s}]=\mathbb{E}[M_{s}\,|\,\mathcal{F}^{M}_{s}]=M_{s}.$ (4.1)
A relationship that we shall use later, between the RAP $(I_{t}^{(1)})$, its noise process $(A_{t}^{(1)})$, and its associated FAM $(M_{t})$ is the following:
@@
If $I_{t}^{(1)}=g_{0}(t)X_{0}+g_{1}(t)X_{1}+A_{t}^{(1)}$ and $M_{t}=\mathbb{E}[X_{1}\,|\,\mathcal{F}_{t}^{I}]$, then
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}]$ (4.2)
@@
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$ $=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
@@
Let $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. We notice that
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$ $=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$
$=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}],$ (4.4)
since $\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,X_{0},X_{1},\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]\,|\,\mathcal{F}^{I}_{s}]$, where we used the fact that $(A_{t}^{(1)})\perp\!\!\!\perp(X_{0},X_{1})$ by Definition 3.1. If $(A_{t}^{(1)})$ is Gauss-Markov, then
$\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{A}_{s}]=\mathbb{E}[A_{t}^{(1)}\,|\,A_{s}^{(1)}]=\mu_{A}(t)+\frac{K_{A}(s,t)}{\sigma_{A}^{2}(s)}\left(A_{s}^{(1)}-\mu_{A}(s)\right).$ (4.5)
@@
[…]
@@
$I_{t}^{(1)}=\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{t}]=\mathbb{E}[A_{t}^{(1)}\,|\,\mathcal{F}^{I}_{t}]+g_{0}(t)X_{0}+g_{1}(t)M_{t},$ (4.7)
@@
$\mathbb{E}[I_{t}^{(1)}\,|\,\mathcal{F}^{I}_{s}]$ $=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
@@
If $(I_{t}^{(1)})$ is $\{T_{0},T_{1}\}$-conditionally-Markov, then $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$. In this case, the dynamics of $(M_{t})$ can be obtained using Bayes’ rule and Itô’s formula under mild assumptions. In the remainder of this section, we assume that the driver of $(I_{t}^{(1)})$ has a density function for $t\in[T_{0},T_{1}]$.
@@
Let $M_{t}=\mathbb{E}[X_{1}\,|\,X_{0},I_{t}^{(1)}]$ be a one-arc FAM restricted to $t\in(T_{0},T_{1})$. Then
$M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)},$ (4.10)
where $f^{I_{t}^{(1)}\,|\,X_{0},X_{1}}$ is the conditional density of $(I_{t}^{(1)})$ given $X_{0}$ and $X_{1}$, and $F^{X_{1}\,|\,X_{0}}$ is the conditional distribution function of $X_{1}$ given $X_{0}$. In particular,
@@
    $M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})f^{X_{1}\,|\,X_{0}}(y)\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})f^{X_{1}\,|\,X_{0}}(y)\,\mathrm{d}y}$ (4.11)
@@
    $M_{t}=\frac{\sum\limits_{y}yf^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathbb{P}[X_{1}=y\,|\,X_{0}]}{\sum\limits_{y}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(I_{t}^{(1)})\mathbb{P}[X_{1}=y\,|\,X_{0}]}.$ (4.12)
@@
$\mathbb{P}[X_{1}\leqslant y\,|\,X_{0},z\leqslant I_{t}^{(1)}\leqslant z+\epsilon]$ $=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}\leqslant y]\mathbb{P}[X_{1}\leqslant y\,|\,X_{0}]}{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0}]}$
$=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}\leqslant y]\mathbb{P}[X_{1}\leqslant y\,|\,X_{0}]}{\int_{\mathbb{R}}\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}=y]\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}$
$=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon,X_{1}\leqslant y\,|\,X_{0}]}{\int_{\mathbb{R}}\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\,|\,X_{0},X_{1}=y]\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)}.$ (4.13)
@@
$F^{X_{1}\,|\,X_{0},I_{t}^{(1)}=z}(y)=\frac{\frac{\,\mathrm{d}}{\,\mathrm{d}z}\mathbb{P}[I_{t}^{(1)}\leqslant z\,|\,X_{0},X_{1}\leqslant y]F^{X_{1}\,|\,X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\,|\,X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\,|\,X_{0}}(y)},$ (4.14)

### v3

@@ 5 Filtered arcade martingales
    $\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})\mathrel{\mathop{\ordinarycolon}}=\{\pi\in\Pi(\mu_{0},\mu_{1},\ldots,\mu_{n})\cap\mathcal{P}_{1}(\mathbb{R}^{n+1})\,|\,(X_{0},X_{1},\ldots,X_{n})\sim\pi\implies\\ \mathbb{E}[X_{n}\,|\,X_{0},\ldots,X_{m}]=X_{m}\text{ for all }m=0,\ldots,n\}$ is the set of martingale couplings of $\{\mu_{0},\mu_{1},\ldots,\mu_{n}\}\subseteq\mathcal{P}(\mathbb{R})$.
We recall that $\mathcal{M}(\mu_{0},\mu_{1},\ldots,\mu_{n})\neq\emptyset$ if and only if $\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{0}(x)\leqslant\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{1}(x)\leqslant\ldots\leqslant\int_{\mathbb{R}}f(x)\,\mathrm{d}\mu_{n}(x)$ for any convex function $f$ on $\mathbb{R}$, i.e, the measures $\mu_{0},\mu_{1},\ldots,\mu_{n}$ are convexly ordered. In that case, we write $\mu_{0}\leqslant_{\mathrm{cx}}\mu_{1}\leqslant_{\mathrm{cx}}\ldots\leqslant_{\mathrm{cx}}\mu_{n}$.
@@ 5.1 The one-arc FAM
Let $X=(X_{0},X_{1})$ be a random vector distributed according to a martingale coupling $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1})$, where $(\mu_{0},\mu_{1})\in\mathcal{P}_{1}(\mathbb{R})\times\mathcal{P}_{1}(\mathbb{R})$ and $\mu_{0}\leqslant_{\mathrm{cx}}\mu_{1}$. It follows that $X_{0}\sim\mu_{0}$, $X_{1}\sim\mu_{1}$, and $\mathbb{E}[X_{1}\mid X_{0}]\overset{a.s.}{=}X_{0}$. Given an $X$-RAP $(I_{t}^{(1)})_{t\in[T_{0},T_{1}]}$ on the partition $\{T_{0},T_{1}\}_{*}$, we aim at constructing an $(\mathcal{F}^{I}_{t})$-martingale $(M_{t})_{t\in[T_{0},T_{1}]}$ such that $M_{T_{0}}\overset{a.s.}{=}X_{0}$ and $M_{T_{1}}\overset{a.s.}{=}X_{1}$. The BHM framework developed in [7] is recovered when $X_{0}=\mathbb{E}[X_{1}]=0$ and $T_{0}=0$.
@@
Given an $X$-RAP $(I_{t}^{(1)})$, a one-arc FAM for $X\sim\pi^{X}$ on $T_{0},T_{1}]$ is a stochastic process defined by $M_{t}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}[X_{1}\mid\mathcal{F}_{t}^{I}]$.
@@
    For $s<t$, $\mathbb{E}[M_{t}\mid\mathcal{F}_{s}^{I}]=M_{s}$ by the tower property of the conditional expectation.
@@
    $M_{T_{0}}=\mathbb{E}[X_{1}\mid X_{0}]\overset{a.s.}{=}X_{0}$, since $\pi^{X}\in\mathcal{M}(\mu_{0},\mu_{1})$ and $I_{T_{0}}^{(1)}=X_{0}$.
@@
    $M_{T_{1}}=\mathbb{E}[X_{1}\mid\mathcal{F}_{T_{1}}^{I}]\overset{a.s.}{=}X_{1}$, since $\sigma(X_{1})\subset\mathcal{F}_{T_{1}}^{I}$ by construction of $(I_{t}^{(1)})$.
Hence, $(M_{t})$ is an $(\mathcal{F}_{t}^{I})$-martingale that interpolates almost surely between $X_{0}$ and $X_{1}$ over the interval $T_{0},T_{1}]$. ∎
@@
$\mathbb{E}[M_{t}\mid\mathcal{F}^{M}_{s}]\overset{a.s.}{=}\mathbb{E}[\mathbb{E}[M_{t}\mid\mathcal{F}^{M}_{s}]\mid\mathcal{F}^{I}_{s}]\overset{a.s.}{=}\mathbb{E}[\mathbb{E}[M_{t}\mid\mathcal{F}^{I}_{s}]\mid\mathcal{F}^{M}_{s}]\overset{a.s.}{=}\mathbb{E}[M_{s}\mid\mathcal{F}^{M}_{s}]\overset{a.s.}{=}M_{s}.$ (5.1)
To assist with the derivation of the outcomes in the remainder of this section, we give a brief overview of the goals and forthcoming results as a guide. First, we recall the martingale representation theorem (MRT), as presented in [27], Theorem 4.2.
Martingale representation theorem, [27]. Let $(W_{t})_{t\geqslant 0}$ be a standard Brownian motion on a filtered probability space $(\Omega,\mathcal{F},\left(\mathcal{F}_{t}\right)_{t\geq 0},\mathbb{P})$ satisfying the usual conditions. [^3] The usual conditions are: 1. Right-continuity of the filtration: $\mathcal{F}_{t}=\bigcap_{s>t}\mathcal{F}_{s}\quad\text{ for all }t\geqslant 0.$ 2. Completeness of the probability space: The $\sigma$-algebra $\mathcal{F}_{0}$ contains all $\mathbb{P}$-null sets of $\mathcal{F}$. That is, if $A\in\mathcal{F}$ and $\mathbb{P}(A)=0$, then $A\in\mathcal{F}_{0}$. 3. Adaptedness of the Brownian motion: The Brownian motion $(W_{t})$ is adapted to the filtration $(\mathcal{F}_{t})$. Then, every square integrable $(\mathcal{F}_{t}$)-martingale $(M_{t})_{t\geq 0}$ can be represented by
$M_{t}=M_{0}+\int_{0}^{t}\phi_{s}\,\mathrm{d}W_{s}$ (5.2)
almost surely, where $(\phi_{t})$ is a predictable process such that $\int_{0}^{T}\phi_{s}^{2}\,\mathrm{d}s<\infty$ almost surely for all $T>0$.
The MRT assumes that the filtration $(\mathcal{F}_{t})$ be generated by a Brownian motion (or at least rich enough to support one). If $(\mathcal{F}_{t})$ is more general or includes information not generated by Brownian motion (e.g., jump processes, discrete-time processes, or other sources of randomness), then:
  • 1.
    A Brownian motion may not exist with respect to the filtration $(\mathcal{F}_{t})$.
  • 2.
    Even if a Brownian motion exists, it may not generate the entire filtration; in this case the MRT would not apply.
The main goals of this section are:
  • 1.
    Show that even though $(\mathcal{F}_{t}^{I})$ is not generated by a Brownian motion, under appropriate conditions, one may write
    $M_{t}\overset{a.s.}{=}X_{0}+\int_{0}^{t}\sigma_{s}\,\mathrm{d}W_{s},$ (5.3)
    where $(W_{t})$ is an $(\mathcal{F}_{t}^{I})$-adapted Brownian motion.
  • 2.
    Identify the processes $(\sigma_{t})$ and $(W_{t})$, i.e., give their explicit, almost sure expressions.
To achieve these goals, we proceed as follows.
  1. 1.
    In Proposition 5.30, we bring to the fore a relationship between $\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]$, $M_{s}$, and $\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{A}_{s}]\mid\mathcal{F}^{I}_{s}]$ for any pair $(s,t)\in[T_{0},T_{1}]^{2}$ such that $s\leqslant t$. This relationship will play a key role in finding the $(\mathcal{F}_{t}^{I})$-adapted Brownian motion $(W_{t})$.
  2. 2.
    In the case where $M_{t}=\mathbb{E}[X_{1}\mid\mathcal{F}_{t}^{I}]=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$ holds, i.e., in the case that $(I_{t}^{(1)})$ is a conditional-Markov process, we apply the Bayes rule to $(M_{t})$ in Proposition 5.31 to obtain an integral representation that depends on the conditional density of $(I_{t}^{(1)})$ given $X_{0}$ and $X_{1}$, and the conditional distribution function of $X_{1}$ given $X_{0}$.
  3. 3.
    We then apply Itô’s formula to the Bayes integral representation in Proposition 5.32. This allows us to obtain the general SDE satisfied by a FAM that is driven by a conditional-Markov RAP, under the appropriate conditions implied by Itô’s formula. It then turns out that the SDE is composed by a $\mathrm{d}t-$term, a $\mathrm{d}I_{t}^{(1)}-$term, and a $\mathrm{d}[I^{(1)}]_{t}\,-$term.
  4. 4.
    We specialise the SDE in two steps. First, we add the assumption that the driver of $(I_{t}^{(1)})$ is Gaussian in Corollary 5.34. This allows us to calculate all the terms in the SDE explicitly. Then, we add the assumption that $(I_{t}^{(1)})$ is a standard RAP, see Corollary 5.35, which allows us to calculate $I^{(1)}]_{t}$ explicitly and further simplify the terms in the SDE. These now only have a $\mathrm{d}t-$term and a $\mathrm{d}I_{t}^{(1)}-$term, in which a common factor appears.
  5. 5.
    Removing that common factor, Proposition 5.37 shows that what is left is an $(\mathcal{F}^{I}_{t})$-adapted Brownian motion, up to rescaling. Thus, the common factor yields the volatility process $(\sigma_{t})$. Hence, one has achieved one’s goal. This plan is rather standard in a non-linear stochastic filtering context, see, e.g., [10].
We now begin with the first step of the described plan, above. Since equalities involving conditional expectations are understood in the almost sure sense, we will omit the a.s. “symbol” in $\overset{a.s.}{=}$" from now on in equalities involving FAMs.
A relationship that we shall use later between the RAP $(I_{t}^{(1)})$, its noise process $(A_{t}^{(1)})$, and its associated FAM $(M_{t})$ is the following:
@@
If $I_{t}^{(1)}=g_{0}(t)X_{0}+g_{1}(t)X_{1}+A_{t}^{(1)}$ and $M_{t}=\mathbb{E}[X_{1}\mid\mathcal{F}_{t}^{I}]$, then
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{A}_{s}]\mid\mathcal{F}^{I}_{s}]$ (5.4)
@@
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]$ $=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
@@
Let $(s,t)\in[T_{0},T_{1}]^{2}$ be such that $s\leqslant t$. We notice that
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]$ $=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]$
$=g_{0}(t)X_{0}+g_{1}(t)M_{s}+\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{A}_{s}]\mid\mathcal{F}^{I}_{s}],$ (5.6)
since $\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\mid X_{0},X_{1},\mathcal{F}^{A}_{s}]\mid\mathcal{F}^{I}_{s}]=\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{A}_{s}]\mid\mathcal{F}^{I}_{s}]$, where we used the fact that $(A_{t}^{(1)})\perp\!\!\!\perp(X_{0},X_{1})$ by Definition 3.16. If $(A_{t}^{(1)})$ is Gauss-Markov, then
$\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{A}_{s}]=\mathbb{E}[A_{t}^{(1)}\mid A_{s}^{(1)}]=\mu_{A}(t)+\frac{K_{A}(s,t)}{\sigma_{A}^{2}(s)}\left(A_{s}^{(1)}-\mu_{A}(s)\right).$ (5.7)
@@
$\mathbb{E}[\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{A}_{s}]\mid\mathcal{F}^{I}_{s}]=\mu_{A}(t)+\frac{K_{A}(s,t)}{\sigma_{A}^{2}(s)}\left(\mathbb{E}[A_{s}^{(1)}\mid\mathcal{F}^{I}_{s}]-\mu_{A}(s)\right).$ (5.8)
@@
$I_{t}^{(1)}=\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{t}]=\mathbb{E}[A_{t}^{(1)}\mid\mathcal{F}^{I}_{t}]+g_{0}(t)X_{0}+g_{1}(t)M_{t},$ (5.9)
@@
$\mathbb{E}[I_{t}^{(1)}\mid\mathcal{F}^{I}_{s}]$ $=\left(g_{0}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{0}(s)\right)X_{0}+\left(g_{1}(t)-\frac{A_{2}(t)}{A_{2}(s)}g_{1}(s)\right)M_{s}$
@@
If $(I_{t}^{(1)})$ is $\{T_{0},T_{1}\}$-conditionally Markov, then $M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$. In this case, the dynamics of $(M_{t})$ can be obtained using Bayes’ rule and Itô’s formula under mild assumptions. In the remainder of this section, we assume that the driver of $(I_{t}^{(1)})$ has a density function for $t\in[T_{0},T_{1}]$. In what follows, when we write $M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$, we assume that $(I_{t}^{(1)})$ is $\{T_{0},T_{1}\}$-conditionally Markov.
@@
Let $M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]$ be a one-arc FAM restricted to $t\in(T_{0},T_{1})$. Then
$M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)},$ (5.12)
where $f^{I_{t}^{(1)}\mid X_{0},X_{1}}$ is the conditional density of $(I_{t}^{(1)})$ given $X_{0}$ and $X_{1}$, and $F^{X_{1}\mid X_{0}}$ is the conditional distribution function of $X_{1}$ given $X_{0}$. In particular,
@@
    $M_{t}=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})f^{X_{1}\mid X_{0}}(y)\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})f^{X_{1}\mid X_{0}}(y)\,\mathrm{d}y}$ (5.13)
@@
    $M_{t}=\frac{\sum\limits_{y}yf^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\mathbb{P}[X_{1}=y\mid X_{0}]}{\sum\limits_{y}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\mathbb{P}[X_{1}=y\mid X_{0}]}.$ (5.14)
@@
$\mathbb{P}[X_{1}\leqslant y\mid X_{0},z\leqslant I_{t}^{(1)}\leqslant z+\epsilon]$ $=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\mid X_{0},X_{1}\leqslant y]\mathbb{P}[X_{1}\leqslant y\mid X_{0}]}{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\mid X_{0}]}$
$=\frac{\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\mid X_{0},X_{1}\leqslant y]\mathbb{P}[X_{1}\leqslant y\mid X_{0}]}{\int_{\mathbb{R}}\mathbb{P}[z\leqslant I_{t}^{(1)}\leqslant z+\epsilon\mid X_{0},X_{1}=y]\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)}.$ (5.15)
@@
$F^{X_{1}\mid X_{0},I_{t}^{(1)}=z}(y)=\frac{\frac{\,\mathrm{d}}{\,\mathrm{d}z}\mathbb{P}[I_{t}^{(1)}\leqslant z\mid X_{0},X_{1}\leqslant y]F^{X_{1}\mid X_{0}}(y)}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(z)\,\mathrm{d}F^{X_{1}\mid X_{0}}(y)},$ (5.16)